### v1

# Stylebreeder : Exploring and Democratizing Artistic Styles through Text-to-Image Models

Abstract
Text-to-image models are becoming increasingly popular, revolutionizing the landscape of digital art creation by enabling highly detailed and creative visual content generation. These models have been widely employed across various domains, particularly in art generation, where they facilitate a broad spectrum of creative expression and democratize access to artistic creation. In this paper, we introduce STYLEBREEDER, a comprehensive dataset of 6.8M images and 1.8M prompts generated by 95K users on Artbreeder, a platform that has emerged as a significant hub for creative exploration with over 13M users. We introduce a series of tasks with this dataset aimed at identifying diverse artistic styles, generating personalized content, and recommending styles based on user interests. By documenting unique, user-generated styles that transcend conventional categories like ‘cyberpunk’ or ‘Picasso,’ we explore the potential for unique, crowd-sourced styles that could provide deep insights into the collective creative psyche of users worldwide. We also evaluate different personalization methods to enhance artistic expression and introduce a style atlas, making these models available in LoRA format for public use. Our research demonstrates the potential of text-to-image diffusion models to uncover and promote unique artistic expressions, further democratizing AI in art and fostering a more diverse and inclusive artistic community. The dataset, code and models are available at https://stylebreeder.github.io under a Public Domain (CC0) license.
1 Introduction
Text-to-image models, such as Denoising Diffusion Models (DDMs) [18] and Latent Diffusion Models (LDMs) [37], are becoming increasingly popular and are revolutionizing the landscape of digital art creation. These models are renowned for their capability to generate high-quality, high-resolution images across diverse domains, enabling the production of highly detailed and creative visual content [57, 8, 20, 58, 12, 31, 24]. Artists worldwide are increasingly leveraging these models, utilizing diverse textual prompts to create artworks spanning myriad styles, thereby democratizing the art creation process and making it more accessible.
In the midst of this technological renaissance, platforms like Artbreeder have surfaced as pivotal hubs for creative exploration. Artbreeder, with its user base exceeding 13 million, facilitates the generation of millions of images that embody a vast spectrum of artistic styles. This surge in user-generated content presents an intriguing question: beyond the conventional styles typically prompted by terms like ‘cyberpunk’ or ‘Picasso,’ what unique, crowd-sourced styles might exist within such a community? These styles, potentially undocumented and uniquely communal, could offer profound insights into the collective creative psyche of users worldwide.
However, existing datasets often fall short in terms of exploring the visual potential of user-generated images. While some works similar to ours, such as Diffusion DB [55], TWIGMA [5] also explore AI-generated content, they either feature a smaller user base, do not provide original text prompts associated with the images, or provide limited insights into artistic styles from a visual perspective. Addressing this inquiry, our paper introduces a comprehensive dataset derived from Artbreeder, providing contributions from 95K unique users, 6.8M images and 1.8 text-prompts generated with a variety of text-to-image diffusion models such as Stable Diffusion [37] or SD-XL [34] (see Fig. 1). We identified and analyzed novel, user-generated artistic styles, uncovering diverse and previously unrecognized creative expressions. Based on the discovered styles, we showcase a variety of tasks including personalization and recommendation. Our contributions are as follows:
We present an extensive dataset from Artbreeder on CC0 license, capturing millions of user-generated images and styles, and sharing with the community to further encourage research in this area.
We cluster images based on their stylistic similarities, helping to map out the landscape of user-generated art.
Utilizing historical data of the users, we showcase a recommendation system that aligns style suggestions with individual preferences, making the exploration of artistic styles more targeted and meaningful.
We release a web-based platform, Style Atlas, providing public access to download pre-trained style LoRAs for personalized content generation, encouraging experimentation and collaborative artistic exploration to expand our understanding of digital creativity and further democratize the creative use of AI.
2 Related Work
We offer a brief overview of artwork datasets, followed by discussions on text-to-image diffusion models, and personalized image generation.
2.1 Artwork Datasets
Traditional artwork datasets (see Supplemental Materials for a comparison) have primarily focused on artwork classification and attribute prediction, however, these datasets often exhibit limitations, like skewed class distributions and unsuitable classes for image synthesis, when employed for artwork synthesis. To address these shortcomings in evaluating artwork synthesis, specialized subsets have been curated to better suit the task. For instance, [50] and [59] derived datasets by scraping WikiArt images. Despite these efforts, such datasets still face many challenges related to variable image quality, imbalanced distributions, and others. ArtBench-10 [27] attempted to rectify these issues by introducing a class-balanced and cleanly-annotated benchmark, however, it only provides 10 classes.
Recent advancements in text-to-image synthesis have spurred the development of AI-generated datasets like DiffusionDB [55] and Midjourney Kaggle [32], which contain millions of image-text pairs generated by models such as Stable Diffusion and Midjourney. These datasets, while groundbreaking, tend to be limited in stylistic diversity and are skewed towards specific user groups, reflecting data collected from constrained environments and short time frames. Researchers have tailored other datasets to investigate certain themes [4, 26, 45, 54, 47, 30, 56], but these are domain specific and lack breadth. TWIGMA [5] captured multiple years of generated images scraped from X (formerly Twitter) but does not include the prompts that were used to generate these images, instead relying on inferred BLIP captions. As shown in Tab. 1, our dataset not only includes the original prompts used to produce images, but also contains images generated by multiple models while encapsulating a long time frame.
| Name | Source | # Images | Year | Original Prompt Included | Multiple Models |
| DiffusionDB | SD Discord | 14,000,000 | Aug 2022 | ✓ | ✗ |
| Midjourney Kaggle | Midjourney | 250,000 | Jun 2022-Jul 2022 | ✓ | ✗ |
| TWIGMA | 800,000 | Jan 2020-Mar 2023 | ✗ | ✓ | |
| STYLEBREEDER (Ours) | Artbreeder | 6,818,217 | Jul 2022-May 2024 | ✓ | ✓ |
2.2 Text-to-image Generative Models
Text-to-image generative models (T2I) [40, 35, 37] have displayed exceptional abilities in synthesizing high-quality images conditioned on textual descriptions. As they improve, these models are becoming increasingly indispensable tools for visually creative tasks. Among these, diffusion-based text-to-image models [19, 2, 46] are prominently used for guided image synthesis [8, 20, 58] and complex image editing [57, 3] applications. Although these models allow for significant control through text, such as directing the color or attributes of styles within generated images, they are unable to consistently preserve the same precise style across new contexts.
2.3 Personalized Generation in Diffusion Models
Personalization techniques have been proposed to enable pre-trained text-to-image models to generate novel concepts based on a small set of images. DreamBooth [38] fine-tunes the full T2I model which yields more detailed and expressive outputs. However, due to the large scale of these models, full fine-tuning is an expensive task that requires substantial amounts of memory. Different methods attempt to work around this challenge for both style and content representations. Custom Diffusion [25] attempts multi-concept learning, but requires expensive joint training and struggles with style disentanglement. Textual Inversion [11] learns a new token embedding to represent a subject or style without altering the original model parameters, while StyleDrop [43] utilizes adapter tuning to only train a subset of weights for style adaptation. Low-Rank Adaptation (LoRA) [21] is a Parameter Efficient Fine-Tuning (PEFT) technique frequently used to fine-tune T2I models to generate images of a desired style. EDLoRA [14] proposes a layerwise embedding and multi-word representation when training a LoRA model.
3 Stylebreeder Dataset
We collected STYLEBREEDER by scraping images from the Artbreeder website. We chose Artbreeder since it is one of the most popular platforms for art generation, supporting a variety of text-to-image models. Additionally, all images on Artbreeder are covered by a CC0 license111https://creativecommons.org/public-domain/cc0, which allows for unrestricted use for any purpose 222https://www.artbreeder.com/terms.pdf. We collected meta data along with the images, which includes text prompt (positive and negative), username, hyperparameters. We provide additional features such as NSFW scores for each images and text prompts (see Fig. 1). We provide a detailed analysis of these features below.
3.1 User Statistics
Our dataset comprises 95,479 unique users, each generating an average of 72.41 images. Figure 2 (a) illustrates the distribution of images per user, showing that the majority of users produced fewer than 1,000 images, although a few power users created a significantly larger number of images. All user IDs in our dataset have been anonymized to ensure the privacy of users.
3.2 Model Statistics
Our dataset represents a wide variety of text-to-image diffusion models, including Stable Diffusion 1.5 (74.9%), SD-XL 1.0 (13.1%), Stable Diffusion 1.3 (8.8%), Stable Diffusion 1.4 (1.3%), Stable Diffusion 1.5-free (1.1%) and ControlNet 1.5 (0.8%). Stable Diffusion 1.5 is the most frequently used model in our dataset followed by SD-XL, which has gained popularity due to its ability to generate high-quality images. The resolution of the images ranges from to based on the model employed. Additionally, our dataset provides details on key hyperparameters like seed, CFG guidance scale—which dictates the extent to which the image generation process adheres to the text prompt—and step size in the diffusion model.
3.3 Text Prompts
We collect both positive and negative text prompts for each image in our dataset. The average prompt length is 60 words, as shown in Fig. 2 (b). Common words in positive prompts, such as ‘painting’, ‘realistic’, and ‘digital’, reveal semantic information about the desired styles of images (see Fig. 2 (c)) while common words in negative prompts like ‘ugly’ and ‘deformed’ indicate the undesired features of the generated images (see Fig. 2 (d)). Furthermore, we observed that users often incorporate artist names in their text prompts to specify desired styles, a common practice employed by generative art community. To quantify this trend, we conducted an analysis using BERT NER [51] to identify unique artist names in the dataset. Our findings highlighted a significant occurrence of artist names, with top mentions including ‘Ilya Kuvshinov’, a Russian illustrator, featured in 208K text prompts, and ‘Akihiko Yoshida’, a Japanese video game artist, appearing in 81K prompts. Given that these artists may not permit the use of their artistic styles, we offer a form on our website allowing artists to opt out, ensuring their styles are not replicated without their consent, as outlined in Section 5.
3.4 NSFW and Toxic Content
We analyze NSFW content in both images and text prompts using state-of-the-art predictors for images [16] and text prompts [15]. Figure 4 (a) shows a comparison of our dataset (STYLEBREEDER) with other popular datasets such as LAION [42], Artbench [27], DiffusionDB [55] and TWIGMA [5]. Most of these datasets, particularly those with AI-generated images like DiffusionDB, TWIGMA, and ours, contain a substantial amount of potentially NSFW content. A similar observation can be made for text-prompts (see Fig. 4 (b)) where we report NSFW, Toxicity, Severe Toxicity, Identity Attack, Obscene, Insult and Threat scores computed with [15]. This trend correlates with recent studies that highlight a significant increase in NSFW content generation by online communities [36, 41]. For instance, potentially harmful text prompts may involve the names of influential politicians, such as ‘Donald Trump’ and ‘Joe Biden’, found in prompts like ‘angry Joe Biden screaming, red-faced, steam coming from ears’ or ‘angry Donald Trump Melania and the judge and police at mcdonalds’. Additionally, our analysis revealed the use of celebrity names in contexts suggesting sexually explicit content that could be construed as nonconsensual pornography. To assist researchers, our dataset includes NSFW text and image scores, along with toxicity-related scores, enabling them to filter these images and determine appropriate thresholds for excluding potentially unsafe data. We also provide a Google form for reporting harmful or inappropriate images and prompts, as outlined in Section 5.
4 Experiments
We have defined three tasks using our dataset: identifying diverse artistic styles through clustering based on stylistic similarities, generating personalized images tailored to individual styles, and recommending styles based on previously generated images of a given user.
4.1 Experimental Setup
We utilized official and Diffusers [33] implementations for Textual Inversion [11], LoRA w/DreamBooth [21, 38], Custom Diffusion [25], and EDLoRA [14]. For each method, we used the same set of seeds and adhered to the parameter settings recommended in their respective publications. Textual Inversion was trained at a resolution of 512x512 for 3000 steps with a learning rate of 5e-4. LoRA with DreamBooth and EDLoRA were both trained at the same resolution (512x512) for 800 steps, but with a learning rate of 1e-4, and LoRA had a rank of 32. Custom Diffusion was trained at a resolution of 512x512 for 250 steps with a learning rate of 1e-5. DINO and CLIP scores in Tab. 2 were computed on 20K images each, and standard deviations are provided. We used 8 NVIDIA L40 GPUs for our experiments. For recommendation tasks, we used 96K images generated by 1434 users with an 80% training and 20% test split, using learning rate of 5e-3 and regularization term 2e-2 and analyze a 5-fold cross validation using the Surprise [48] library.
4.2 Discovering Diverse Artistic Styles
The rising popularity of generated art has led to a vast array of user-generated content, showcasing a diverse spectrum of artistic styles. However, a key challenge remains: how can we effectively discover and categorize these styles? Our goal is to cluster user-generated images into stylistically similar groups to uncover unique styles. Formally, our dataset, denoted as , consists of images. We employ a pre-trained text-to-image model , parameterized by weights and a token embedding space . Firstly, we convert the images into a set of style embeddings using a state-of-the-art feature extractor , specifically CSD [44], which uses a Vision Transformer (ViT) [9] backbone to map images into a -dimensional vector space representing their style descriptors. CSD has shown superior performance in style-matching tasks across datasets like DomainNet, WikiArt, and LAION-Styles, outperforming models like DINO which focus more on semantic content. This conversion results in a set of embeddings , where each image, , is embedded in a high-dimensional semantic embedding space. These embeddings are then clustered into groups using the K-Means++ [1] algorithm, which utilizes cosine similarity to ensure cohesion within clusters. Figure 4 (c) presents a 2D projection of the CSD embedding space using t-SNE [17]. The visualization reveals that some clusters are closely grouped, reflecting stylistic similarities, while others are distinctly isolated, demonstrating significant stylistic variations.
4.3 Personalized Image Generation Based on Style
Personalized image generation based on style is a crucial task since it allows users to create unique, customized visuals that resonate with their specific aesthetic preferences. We picked four state-of-the-art personalization methods, namely, Textual Inversion [11], LoRA w/DreamBooth [21, 38], Custom Diffusion [25], and EDLoRA [14], to provide a benchmark analysis on the discovered styles. For this task, we randomly selected 40 clusters, and generated 50 images for 10 text-prompts, totaling to 500 images per cluster for each method. To provide a benchmark on these models, we calculate several metrics: CLIP-T (image-text similarity between generated images and text prompts), CLIP-I (image-image similarity between clusters and generated images), and DINO (image-image similarity between clusters and generated images). Details of these metrics can be found in Tab. 2, which demonstrates that EDLoRA exhibits superior performance in personalized image generation due to its embedding-based approach. In addition to the quantitative analysis, qualitative examples are provided in Fig. 6.
| Textual Inversion | LoRA w/DreamBooth | EDLoRA | Custom-Diffusion | ||
| CLIP-I | Avg. | 0.6869 0.10 | 0.6299 0.11 | 0.6957 0.11 | 0.5917 0.12 |
| Min. | 0.6166 0.10 | 0.5654 0.11 | 0.6214 0.11 | 0.5324 0.11 | |
| Max. | 0.7428 0.10 | 0.6831 0.11 | 0.7521 0.12 | 0.6440 0.12 | |
| CLIP-T | Avg. | 0.1857 0.02 | 0.1896 0.02 | 0.1822 0.01 | 0.1809 0.02 |
| Min. | 0.1555 0.02 | 0.1573 0.02 | 0.1527 0.01 | 0.1486 0.02 | |
| Max. | 0.2392 0.03 | 0.2663 0.03 | 0.2389 0.03 | 0.2585 0.03 | |
| DINO | Avg. | 0.3801 0.15 | 0.2668 0.17 | 0.4125 0.18 | 0.2546 0.17 |
| Min. | 0.2581 0.13 | 0.1682 0.14 | 0.2790 0.15 | 0.1634 0.14 | |
| Max. | 0.4838 0.17 | 0.3585 0.19 | 0.5246 0.19 | 0.3402 0.19 |
4.4 Style-based Recommendation
The sheer volume of styles generated by users presents a significant challenge in navigating the landscape of artistic options available. To address this, we showcase a recommendation system that suggests top styles to users based on their previously generated images. This personalized approach is crucial as it helps users discover styles that align with their tastes and past preferences, simplifying their search among a vast array of choices. We formulate our task as a matrix-factorization based recommendation which involves a set of items, where users rate items they have interacted with, thus creating a matrix of user-item ratings. In the context of our problem, users are the creators that generate images, and items are the clusters in which generated images are assigned. We calculate for each user a vector where represents the proportion of images that user has generated within a specific cluster such that . These vectors create a matrix where entries denotes the rating for user for cluster . We employ the SVD algorithm [49, 10] to obtain a prediction for all where is the global average rating. and are the user and item bias terms, respectively. and are the latent factor vectors for item and user , respectively. We minimize the regularized squared error loss and update parameters using Stochastic Gradient Descent [13]. We assess the MAE and RMSE of the predicted ratings against the ground-truth ratings obtaining a mean RMSE of 0.1425 and standard deviation 0.0017 along with a mean MAE of 0.082 and standard deviation 0.001 across all folds. Fig. 5 (d) depicts an example of recommendations for a user based on previously generated styles.
4.5 Style Atlas for Democratizing Artistic Styles
Since LoRA has become a popular tool for lightweight concept tuning within the community [39], we are including 100 style LoRAs in our style atlas. Fig. 5 (c) displays a screenshot of our platform where users can browse and download LoRA models for styles they find appealing. Additionally, we provide a notebook that enables users to load these downloaded LoRA models and generate personalized images. For further details, please refer to the Supplementary Material.
5 Limitations and Societal Impact
While our work significantly advances the integration of AI in creative processes, it also presents certain limitations and societal impacts that warrant careful consideration. One of the primary limitations lies in the potential for over-reliance on technology in artistic creation, which could diminish the value and perception of human-driven artistry and creativity. Additionally, the use of AI in art generation raises concerns about copyright and originality, especially when styles closely mimic those of existing artists without clear attribution. From a societal perspective, while our tools aim to democratize art creation, there exists a risk of reinforcing existing biases present in the training data, which could skew the diversity and representation in generated artworks. Moreover, as these technologies become more accessible, there is a potential for misuse, such as creating deceptive images or deepfakes, which could have broader implications for trust and authenticity in digital media. Acknowledging and addressing these challenges is crucial as we continue to explore the intersection of AI and art.
6 Conclusion
This paper has demonstrated the significant potential of text-to-image diffusion models to explore and catalog the rich tapestry of user-generated artistic styles on the Artbreeder platform. We have successfully identified unique, previously uncharted artistic expressions and demonstrated their application in generating personalized images. Additionally, we demonstrated that a personalized recommendation system enhances user engagement by aligning suggested styles with individual preferences. We also release of an atlas of styles democratizes the access to these innovations, allowing users to experiment with and adopt new artistic expressions. This work not only advances the technological capabilities of AI in the field of art but also contributes to a more inclusive and diverse artistic community. Our dataset offers numerous avenues for further exploration, such as refining the effectiveness of text prompts through iterative adjustments, studying trends in art over time, recommending styles based on both image and textual content, developing a search system for the generated images, and exploring explainable creativity [29].
Appendix A Author Statement
We, the authors, bear all responsibility in case of violation of rights, etc., and confirm that STYLEBREEDER is available under a Public Domain (CC0) license.
Appendix B Dataset Hosting
Please visit our Hugging Face page for viewing and downloading the dataset. Additional information regarding STYLEBREEDER can be found at https://stylebreeder.github.io. We confirm that as authors of STYLEBREEDER, we will ensure necessary maintenance related to the dataset.
Appendix C Croissant Metadata
The croissant metadata can be found at https://huggingface.co/api/datasets/stylebreeder/stylebreeder/croissant
Appendix D Style Atlas
Our Style Atlas of 100 LoRAs open for download is available at https://stylebreeder.github.io/atlas/ and Fig. 7 showcases the website.
Appendix E More Comparisons with Other Datasets
Tab 3 features several prominent image synthesis benchmarks. Certain benchmarks focus on single-category image datasets for unconditional image synthesis evaluation, such face datasets like CelebA [28] and FFHQ [23] (human faces) and MetFaces [22] (face artworks). There are also multi-class datasets like STL-10 [6] and ImageNet [7], but these datasets predominantly consist of photographic images with a limited representation of artwork. ArtBench-10 [27] is primarily composed of artwork, however, with only 10 classes, it heavily skews toward artwork of North American, European, and East Asian origin and fails to encompass digital and modern art. DiffusionDB [55] and TWIGMA [5] present AI-generated art datasets, but suffer from a lack of model diversity and original generation prompts, respectively.
| Name | Min Resolution | Max Resolution | # Images | Domain |
| MetFaces [22] | (1024, 1024) | (1024, 1024) | 1,336 | Faces (art) |
| STL-10 [6] | (96, 96) | (96, 96) | 13,000 | Objects |
| ArtBench-10 [27] | (32, 32) | (10629, 7437) | 60,000 | Artworks |
| FFHQ [23] | (256, 256) | (1024, 1024) | 70,000 | Faces (Flickr) |
| CelebA [28] | (64, 64) | (1024, 1024) | 202,599 | Faces (celebrities) |
| TWIGMA [5] | 5122 | Varied | 800,000 | AI artworks |
| ImageNet [7] | (32, 32) | (256, 256) | 1,431,167 | Objects |
| DiffusionDB [55] | (512, 512) | Varied | 14,000,000 | AI artworks |
| Ours | (512, 512) | (1280, 986) | 6,818,217 | AI artworks |
Appendix F Artist names
On platforms such as Artbreeder, it is common to use artist names in text-prompts to indicate a particular style. We used a popular NER library [52] to extract artist names from the text prompts. Top 20 artist names are provided in Table 4. We provided an additional column in our dataset to indicate potential artist names in the text-prompts. Moreover, we provide a Google form on our website if any artist whose name is used would like to opt-out from our dataset.
| Artist Name | Occurrence |
| Tom Bagshaw | 812355 |
| Stanley Artgerm | 547422 |
| Greg Rutkowski | 521464 |
| Artgerm | 439421 |
| Daniel F Gerhartz | 430276 |
| WLOP | 389215 |
| Charlie Bowater | 356740 |
| Atey Ghailan | 351338 |
| Andrew Atroshenko | 336390 |
| Rossdraws | 289541 |
| Edouard Bisson | 229375 |
| Alphonse Mucha | 211639 |
| Ilya Kuvshinov | 206632 |
| Mike Mignola | 196128 |
| Pino Daeni | 123757 |
| Krenz Cushart | 120184 |
| Ismail Inceoglu | 107547 |
| Victoria Justice | 104879 |
| Luis Royo | 100998 |
| Guweiz | 99543 |
| Adriana Lima | 84967 |
Appendix G Datasheet
[linecolor=violet]
MOTIVATION
For what purpose was the dataset created?
Was there a specific task in mind? Was there
a specific gap that needed to be filled? Please provide a description.
As artists worldwide are increasingly leveraging text-to-image diffusion models to create artworks spanning diverse set of styles, there is a pressing need uncover and promote unique artistic expressions. This will enable users to discover unique artistic styles, generate personalized content using those styles, and get recommendations about styles they would be interested. STYLEBREEDER project was created to address these questions, and to the best of our knowledge, has the largest userbase (95K) spanning to 6.8M images with 1.8M unique text prompts.
Who created this dataset (e.g., which team, research group) and on behalf
of which entity (e.g., company, institution, organization)?
This dataset is collected by Matthew Zheng from Virginia Tech and Enis Simsar from ETH Zurich and Joel Simon from Artbreeder.
What support was needed to make this dataset?
(e.g.who funded the creation of the dataset? If there is an associated
grant, provide the name of the grantor and the grant name and number, or if
it was supported by a company or government agency, give those details.)
No specific funding was used.
Any other comments?
None
[linecolor=violet]
COMPOSITION
What do the instances that comprise the dataset represent (e.g., documents,
photos, people, countries)?
Are there multiple types of instances (e.g., movies, users, and ratings;
people and interactions between them; nodes and edges)? Please provide a
description.
Each instance consists of an image generated by a text-to-image diffusion model (such as SD-XL or ControlNet) and the original text prompt used for generating the image, as well as original metadata and expanded features created by us. The metadata originally collected from the Artbreeder website are: Positive Prompt, Negative Prompt, anonymized ImageID, anonymized UserID, Timestamp, and Image Size (height, width), and model-related hyperparameters including Model Type, Seed, Step, and CFG Scale. We also offer further metadata like Cluster ID, along with scores for Prompt NSFW, Image NSFW, Toxicity, Insult, Threat, Identity Attack computed with [15, 16]. We also provide potential artist names that are used in the text prompt such as “Ilya Kuvshinov”.
How many instances are there in total (of each type, if appropriate)?
6,818,217 images created by 95,479 unique users, with 1.8M unique text-prompts.
Does the dataset contain all possible instances or is it a sample (not
necessarily random) of instances from a larger set?
If the dataset is a sample, then what is the larger set? Is the sample
representative of the larger set (e.g., geographic coverage)? If so, please
describe how this representativeness was validated/verified. If it is not
representative of the larger set, please describe why not (e.g., to cover a
more diverse range of instances, because instances were withheld or
unavailable).
No tests were run in order to determine representativeness of the dataset. However, our dataset consists of all publicly generated images on Artbreeder Platform’s Collage tool, from July 2022-May 2024. Out of this dataset, we provide a 2M version which includes all data as well as the original images, hosted on a publicly available platform: https://huggingface.co/datasets/stylebreeder/stylebreeder. We provide all features except images for the full 6.8M dataset on the same website, but provide a script to scrape the 6.8M images from Artbreeder platform. If requested by the reviewer committee at NeurIPS, we will try our best to find a platform that can accommodate the full image data.
What data does each instance consist of?
“Raw” data (e.g., unprocessed text or images) or features? In either case,
please provide a description.
Each instance contains a unique image ID, corresponding meta and expanded features including Positive Prompt, Negative Prompt, anonymized UserID, Timestamp, and Image Size (height, width), and model-related hyperparameters including Model Type, Seed, Step, and CFG Scale, Cluster ID, Prompt NSFW, Image NSFW, Toxicity, Insult, Threat, Identity Attack scores, Artist names. We also provide the original image files along with 2M subset https://huggingface.co/datasets/stylebreeder/stylebreeder, and a script to download 6.8M version.
Is there a label or target associated with each instance?
If so, please provide a description.
For each image, we provide an extended set of labels including text prompt, model parameters, and other features as discussed above.
Is any information missing from individual instances?
If so, please provide a description, explaining why this information is
missing (e.g., because it was unavailable). This does not include
intentionally removed information, but might include, e.g., redacted text.
There is no missing data.
Are relationships between individual instances made explicit (e.g., users’
movie ratings, social network links)?
If so, please describe how these relationships are made explicit.
Each image is connected to an anonymized user ID. By filtering the dataset by user ID, one can infer relationship between different samples, where those samples are created by the same user.
Are there recommended data splits (e.g., training, development/validation,
testing)?
If so, please provide a description of these splits, explaining the
rationale behind them.
We do not provide any recommended splits.
Are there any errors, sources of noise, or redundancies in the dataset?
If so, please provide a description.
No, all images and related metadata are collected directly from Artbreeder platform.
Is the dataset self-contained, or does it link to or otherwise rely on
external resources (e.g., websites, tweets, other datasets)?
If it links to or relies on external resources, a) are there guarantees
that they will exist, and remain constant, over time; b) are there official
archival versions of the complete dataset (i.e., including the external
resources as they existed at the time the dataset was created); c) are
there any restrictions (e.g., licenses, fees) associated with any of the
external resources that might apply to a future user? Please provide
descriptions of all external resources and any restrictions associated with
them, as well as links or other access points, as appropriate.
We provide a 2M version which includes all data as well as the original images, hosted on a publicly available platform: https://huggingface.co/datasets/stylebreeder/stylebreeder. We provide all features except images for the full 6.8M dataset on the same website, but provide a script to scrape the 6.8M images from Artbreeder platform. If requested by the reviewer committee at NeurIPS, we will try our best to find a platform that can accommodate the full image data.
Does the dataset contain data that might be considered confidential (e.g.,
data that is protected by legal privilege or by doctor-patient
confidentiality, data that includes the content of individuals’ non-public
communications)?
If so, please provide a description.
The dataset is collected from user’s publicly available content, however since there is no control over what type of text prompts users can input, it is possible that some confidential information might be included in the text prompts, however we expect such cases to be rare.
Does the dataset contain data that, if viewed directly, might be offensive,
insulting, threatening, or might otherwise cause anxiety?
If so, please describe why.
Similar to other datasets with AI-generated images like DiffusionDB, TWIGMA, ours contain a substantial amount of potentially NSFW content both in terms of images and text prompts. This observation correlates with recent studies that highlight a significant increase in NSFW content generation by online communities [36, 41]. We provide an analysis on NSFW and other features such as Toxicity, Severe Toxicity, Identity Attack, Obscene, Insult and Threat in Fig 3 and Section 3.4 in the main paper. Moreover, to help researchers to filter out potentially unsafe content we provide these scores in our dataset.
Does the dataset relate to people?
If not, you may skip the remaining questions in this section.
We share anonymized user IDs and text prompts may have celebrity or artist names in the text prompts.
Does the dataset identify any subpopulations (e.g., by age, gender)?
If so, please describe how these subpopulations are identified and
provide a description of their respective distributions within the dataset.
No.
Is it possible to identify individuals (i.e., one or more natural persons),
either directly or indirectly (i.e., in combination with other data) from
the dataset?
If so, please describe how.
Text prompts may have celebrity or artist names in the text prompts. We provide a Google form for reporting harmful or inappropriate images and prompts, as well as allowing artists to opt-out in case their names are used in the text prompts.
Does the dataset contain data that might be considered sensitive in any way
(e.g., data that reveals racial or ethnic origins, sexual orientations,
religious beliefs, political opinions or union memberships, or locations;
financial or health data; biometric or genetic data; forms of government
identification, such as social security numbers; criminal history)?
If so, please provide a description.
We do not collect any user-related data except anonymized user ID.
Any other comments?
No.
[linecolor=violet]
COLLECTION
How was the data associated with each instance acquired?
Was the data directly observable (e.g., raw text, movie ratings),
reported by subjects (e.g., survey responses), or indirectly
inferred/derived from other data (e.g., part-of-speech tags, model-based
guesses for age or language)? If data was reported by subjects or
indirectly inferred/derived from other data, was the data
validated/verified? If so, please describe how.
The images, prompt and other metadata including including Positive Prompt, Negative Prompt, anonymized ImageID, anonymized UserID, Timestamp, and Image Size (height, width), and model-related hyperparameters including Model Type, Seed, Step, and CFG Scale are collected from Artbreeder website. Moreover we derived additional features such as Cluster ID, Prompt NSFW, Image NSFW, Toxicity, Insult, Threat, Identity Attack scores, Artist names [15, 16, 53].
Over what timeframe was the data collected?
Does this timeframe match the creation timeframe of the data associated
with the instances (e.g., recent crawl of old news articles)? If not,
please describe the timeframe in which the data associated with the
instances was created. Finally, list when the dataset was first published.
All data was collected from April-May 2024, The images generated between July 2022-May 2024. The dataset first published with this NeurIPS submission, June 10, 2024.
What mechanisms or procedures were used to collect the data (e.g., hardware
apparatus or sensor, manual human curation, software program, software
API)?
How were these mechanisms or procedures validated?
The image IDs and original metadata is directly pulled from the database by one of the authors Joel Simon, owner of Artbreeder platform. A Python script was used to scrape the images, and infer additional features such as Cluster ID, or NSFW scores.
What was the resource cost of collecting the data?
(e.g. what were the required computational resources, and the associated
financial costs, and energy consumption - estimate the carbon footprint.
See Strubell et al. for approaches in this area.)
The metadata and image IDs were provided directly by Artbreeder via their database. Collecting the corresponding images took a total of 78 hours, estimated bandwidth Costs: $315, storage Costs: $80.5 and Energy Costs: $3.9 (assuming an electricity cost of $0.10 per kWh to run a server for 78 hours).
If the dataset is a sample from a larger set, what was the sampling
strategy (e.g., deterministic, probabilistic with specific sampling
probabilities)?
We collected all data from Artbreeder’s Collage tool from July 2022-May 2024, so this is not a subset.
Who was involved in the data collection process (e.g., students,
crowdworkers, contractors) and how were they compensated (e.g., how much
were crowdworkers paid)?
Students collected the data as a part of their graduate research assistantship by Virginia Tech or ETH Zurich.
Were any ethical review processes conducted (e.g., by an institutional
review board)?
If so, please provide a description of these review processes, including
the outcomes, as well as a link or other access point to any supporting
documentation.
No ethical review process was conducted.
Does the dataset relate to people?
If not, you may skip the remainder of the questions in this section.
Our dataset relates to people in 2 ways: 1) We share anonymized user IDs for users who generate the images. 2) Some text prompts may have celebrity or artist names in the prompt.
Did you collect the data from the individuals in question directly, or
obtain it via third parties or other sources (e.g., websites)?
We collected the data from the Artbreeder website.
Were the individuals in question notified about the data collection?
If so, please describe (or show with screenshots or other information) how
notice was provided, and provide a link or other access point to, or
otherwise reproduce, the exact language of the notification itself.
No, by Artbreeder’s contract every user accepts the agreement that their data is licensed as CC0 (Public Domain). Exact language can be seen from: https://www.artbreeder.com/terms.pdf.
Did the individuals in question consent to the collection and use of their
data?
If so, please describe (or show with screenshots or other information) how
consent was requested and provided, and provide a link or other access
point to, or otherwise reproduce, the exact language to which the
individuals consented.
Artbreeder’s contract every user accepts the agreement that their data is licensed as CC0 (Public Domain), hence we did not ask for additional consent.
If consent was obtained, were the consenting individuals provided with a
mechanism to revoke their consent in the future or for certain uses?
If so, please provide a description, as well as a link or other access
point to the mechanism (if appropriate)
Users will have the option to report any harmful, unsafe and sensitive content via our Google form on our website: http://stylebreeder.github.io. Also, artists whose names are used in the text prompts has the option to opt-out via the same form.
Has an analysis of the potential impact of the dataset and its use on data
subjects (e.g., a data protection impact analysis) been conducted?
If so, please provide a description of this analysis, including the
outcomes, as well as a link or other access point to any supporting
documentation.
No data protection impact analysis has been conducted.
Any other comments?
No.
[linecolor=violet]
PREPROCESSING / CLEANING / LABELING
Was any preprocessing/cleaning/labeling of the data
done(e.g.,discretization or bucketing, tokenization, part-of-speech
tagging, SIFT feature extraction, removal of instances, processing of
missing values)?
If so, please provide a description. If not, you may skip the remainder of
the questions in this section.
We used a library to infer various scores such as NSFW, Toxicity, Insult, Threat, Identity Attack scores using Detoxify library [15]. Image NSFW scores are computed using NSFW-Detector library [16]. To infer artist names, we used Named Entity Recognition via [53].
Was the “raw” data saved in addition to the preprocessed/cleaned/labeled
data (e.g., to support unanticipated future uses)?
If so, please provide a link or other access point to the “raw” data.
Yes, raw data is saved.
Is the software used to preprocess/clean/label the instances available?
If so, please provide a link or other access point.
All of our data collection and preprocessing code is available on our website: http://stylebreeder.github.io
Any other comments?
No.
[linecolor=violet]
USES
Has the dataset been used for any tasks already?
If so, please provide a description.
In our paper, we showcased a variety of tasks; 1) discovering diverse styles using clustering based on style representations, 2) personalized content generation on these styles with LoRA and similar methods 3) style recommendation.
Is there a repository that links to any or all papers or systems that use the dataset?
If so, please provide a link or other access point.
This is the first paper using this dataset, however if there are other papers or systems use this dataset in the future, we will include a link in our website.
What (other) tasks could the dataset be used for?
As discussed in our paper, our dataset offers numerous avenues for further exploration, such as refining the effectiveness of text prompts through iterative adjustments, studying trends in art over time, recommending styles based on both image and textual content, developing a search system for the generated images, and exploring explainable creativity.
Is there anything about the composition of the dataset or the way it was
collected and preprocessed/cleaned/labeled that might impact future uses?
For example, is there anything that a future user might need to know to
avoid uses that could result in unfair treatment of individuals or groups
(e.g., stereotyping, quality of service issues) or other undesirable harms
(e.g., financial harms, legal risks) If so, please provide a description.
Is there anything a future user could do to mitigate these undesirable
harms?
Since we collected full data as opposed to a subset, we do not expect any unfair treatment of individuals or groups. Moreover, since the dataset is already in public domain, we do not expect any legal or financial risks. However, as discussed in our paper, some of the images and text prompts reflect NSFW content, and researchers should use the associated scores provided by us to filter out unsafe data based on their tasks.
Are there tasks for which the dataset should not be used?
If so, please provide a description.
All tasks utilize this dataset should follow CC0 Public domain licensing rules.
Any other comments?
No.
[linecolor=violet]
DISTRIBUTION
Will the dataset be distributed to third parties outside of the entity
(e.g., company, institution, organization) on behalf of which the dataset
was created?
If so, please provide a description.
Yes, the dataset is available on the internet.
How will the dataset will be distributed (e.g., tarball on website, API,
GitHub)?
Does the dataset have a digital object identifier (DOI)?
We share the dataset, code and other information on our website; http://stylebreeder.github.io.
When will the dataset be distributed?
The dataset is released on June 12, 2024.
Will the dataset be distributed under a copyright or other intellectual
property (IP) license, and/or under applicable terms of use (ToU)?
If so, please describe this license and/or ToU, and provide a link or other
access point to, or otherwise reproduce, any relevant licensing terms or
ToU, as well as any fees associated with these restrictions.
All images generated on Artbreeder platform are under the CC0 Public Domain License, therefore our images, text prompts and other metadata follows the same rules.
Have any third parties imposed IP-based or other restrictions on the data
associated with the instances?
If so, please describe these restrictions, and provide a link or other
access point to, or otherwise reproduce, any relevant licensing terms, as
well as any fees associated with these restrictions.
All images generated on Artbreeder platform are under the CC0 Public Domain License.
Do any export controls or other regulatory restrictions apply to the
dataset or to individual instances?
If so, please describe these restrictions, and provide a link or other
access point to, or otherwise reproduce, any supporting documentation.
No.
Any other comments?
No.
[linecolor=violet]
MAINTENANCE
Who is supporting/hosting/maintaining the dataset?
Authors of this paper are responsible for supporting, hosting and maintaining the dataset. We also provide a Google form for users to report any potential issue about the dataset.
How can the owner/curator/manager of the dataset be contacted (e.g., email
address)?
The contact information is listed on our website: http://stylebreeder.github.io. We also provide a Google form for users to report any potential issue about the dataset.
Is there an erratum?
If so, please provide a link or other access point.
There is no erratum with our initial release, but we will update our website if we have any errors in the future.
Will the dataset be updated (e.g., to correct labeling errors, add new
instances, delete instances)?
If so, please describe how often, by whom, and how updates will be
communicated to users (e.g., mailing list, GitHub)?
Authors of this paper will monitor the Google form where users can report any potential issues. We will update our datasets biweekly, and announce the updates on our website: http://stylebreeder.github.io.
If the dataset relates to people, are there applicable limits on the
retention of the data associated with the instances (e.g., were individuals
in question told that their data would be retained for a fixed period of
time and then deleted)?
If so, please describe these limits and explain how they will be enforced.
Users can contact us from the Google form and report any potentially harmful or sensitive information for removal.
Will older versions of the dataset continue to be
supported/hosted/maintained?
If so, please describe how. If not, please describe how its obsolescence
will be communicated to users.
Older versions of the dataset will be hosted.
If others want to extend/augment/build on/contribute to the dataset, is
there a mechanism for them to do so?
If so, please provide a description. Will these contributions be
validated/verified? If so, please describe how. If not, why not? Is there a
process for communicating/distributing these contributions to other users?
If so, please provide a description.
Other researchers are welcome to extend or contribute to our dataset, they can either send a pull request to our github repo (http://github.com/stylebreeder) or contact authors on Google form.
Any other comments?
No.